# From loop quantum gravity to cosmology: the 2-vertex model

### v2

@@ Abstract
We study the notion of volume and its dynamics in the loop-quantum-gravity truncation known as the 2-vertex model. We also show that its U($N$)-symmetry reduction provides the old effective dynamics of loop quantum cosmology with an arbitrary perfect barotropic fluid content. A suitable modification of the Poisson bracket structure of the U($N$)-symmetric model leads to the LQC improved dynamics.
@@ I Introduction
The study of simple (truncated) models within the full theory has proved to be extremely useful. The truncation to a fixed graph is described by the corresponding holonomy-flux phase space on the graph and its quantization leads to the spin networks with support on that graph. The simplest non-trivial graphs are the 2-vertex, $N$-edge graphs (the so-called 2-vertex model, for brevity) [13, 14, 15, 16, 17, 18]. The 2-vertex model can be interpreted as two polyhedra with identical areas but, in general, with different volume. The classical dynamics of these polyhedra has been studied analytically in a symmetry-reduced sector (by a global $\text{U}(N)$ symmetry) [14] and, more recently, numerically in the general case [17]. In both situations a cosmological behavior is observed, presenting oscillatory and divergent regimes that avoid any singularities. Besides, in [18] it was shown that, in the low-curvature regime, the $\text{U}(N)$-reduced 2-vertex dynamics is related to the classical dynamics of a Friedmann-Lemaître-Robertson-Walker (FLRW) universe with a cosmological constant.
We will describe the 2-vertex model using the spinorial formalism for LQG [19, 14, 20, 21], a description of the phase space of each edge using two spinors, one for each vertex of the edge. These spinors satisfy two constraints: the matching constraint—which ensures that there is only one SU(2) irreducible representation (one spin) on each edge—and the closure constraint—which ensures the SU(2) invariance of the theory. The spinorial formalism allows for a simple description of the model in terms of $2N$ spinors that, upon quantization, recovers the Hilbert space of LQG on the graph [19]. Therefore, we can study the classical dynamics of the model using the spinorial formalism in order to gain insight into full LQG, as it will give us a first contribution to the quantum dynamics of states peaked on classical trajectories. Moreover, using Minkowski’s theorem for convex polyhedra [22], the spinors provide a parametrization of the framework described by the twisted geometries [23, 24, 21]. In this framework, each vertex of the graph is associated with a polyhedron and the edges indicate adjacent faces with the same area but, in general, different shapes. In this way, the classical counterpart (parametrized by the spinors) associated with a spin network gives us an unconventional discretization of space in which the polyhedra do not fit, i.e., they are twisted with respect to each other.
@@
The $\text{U}(N)$-symmetric sector of the 2-vertex model contains just one degree of freedom corresponding to the twist angle and its canonical conjugate momentum, i.e., the total area of the polyhedra. Comparing the evolution of the quadrupole approximation to the volume and the total area, we find that they precisely follow the evolution one expects from a homogeneous and isotropic evolution. We also find that a generalization of the 2-vertex Hamiltonian constraint in the form of the addition of a function of the area (that preserves the symmetries of the reduced model) effectively takes into account matter content and, furthermore, fully determines the equation of state.
Finally, we show that the 2-vertex model is indeed an LQG truncation that effectively describes the LQC old dynamics [28, 29] with arbitrary barotropic perfect fluids. Furthermore, a suitable modification of the Poisson bracket structure of the U($N$)-symmetric model leads to the LQC improved dynamics [9].
The paper is structured as follows. In Sec. II we introduce the spinorial formalism and the twisted geometries. We will use both of them in Sec. III to describe the 2-vertex model and its dynamics. We will also use the notion of quadrupole moment of a discrete surface to study analytically the volume and its evolution. In Sec. IV we introduce the global $\text{U}(N)$ symmetry and study the dynamics on the reduced sector. In Sec. V, we show that the 2-vertex model directly provides the FRLW cosmology and the LQC effective dynamics. We finally summarize and conclude in Sec. VI.
Index notation: $a,b\ldots=1,2,3$ are spatial indices; $i,j\ldots=1,2...,N$ denote the edges of the 2-vertex graph; $I,J\ldots=1,2,3$ are $\mathfrak{su}(2)$ indices; and $A,B=0,1$ are spinorial indices.
@@ II.1 Spinorial formalism for LQG
This parametrization of $T^{*}\mathrm{SU(2)}$ in terms of two spinors is the classical analogue to Schwinger’s representation of SU(2) by two decoupled harmonic oscillators  [30]. According to this representation, by canonically quantizing the spinors with the commutation relations (3), we will obtain two harmonic oscillators on each edge, describing the SU(2) irreducible representation associated with such edge [30, 15, 19].
@@
[…]
$\vec{X}^{s,t}_{i}\!\mathrel{\mathop{\ordinarycolon}}=\!\frac{1}{2}\!\bra{z_{i}^{s,t}}\vec{\sigma}\ket{z_{i}^{s,t}},\quad\!\!X^{s,t}_{i}\mathrel{\mathop{\ordinarycolon}}=|\vec{X}^{s,t}_{i}|\!=\!\frac{1}{2}\!\braket{z_{i}^{s,t}}{z_{i}^{s,t}}.$ (6)
@@
The Barbero-Immirzi parameter $\beta$ divides the flux and (non-trivially) modifies the phases of the holonomy, in such a way that the Poisson bracket is invariant under the transformation (4) as should in LQG. The action of $g_{i}$ associates the (normalized) spinors on one vertex with the dual (normalized) spinors on the other vertex:
@@ II.2 Twisted geometries
The closure constraint (10) on non-coplanar vectors is precisely the condition required by Minkowski’s theorem [22]. Therefore, these vectors define a unique polyhedron at each vertex, whose faces are orthogonal to the vectors and its areas equal their norms. The faces associated with the same edge but different vertex are understood as adjacent, and the matching constraint (9) causes both faces on each edge to have the same area. However, there is no restriction on the shape of the faces. The faces associated with the same edge do not necessarily have to fit together, giving a notion of twist between adjacent faces. The twist on each edge is parameterized by the twist angle $\xi_{i}\in[-\pi,\pi]$. Consequently, a graph will describe a discretization of space, in which the polyhedra that share an edge will have an adjacent face with the same area but different shapes. For this reason this formalism is known as twisted geometries [23, 24]. According to this formalism, the phase space corresponding to each edge $i$ connecting two vertices is described by $(\hat{X}^{s}_{i},\hat{X}^{t}_{i},X_{i},\xi_{i})$, where $\hat{X}^{s}_{i}$ and $\hat{X}^{t}_{i}$ are the normal vectors to the source and target faces corresponding to the edge $i$; $X_{i}\mathrel{\mathop{\ordinarycolon}}=X_{i}^{s}=X_{i}^{t}$ is the area of the faces associated with the edge $i$; and $\xi_{i}$ is the twist angle of that edge. See Fig. (1) for a graphical representation of the polyhedra associated with a given graph.
@@
III 2-vertex model
We will now apply the formalisms explained in the previous section to describe the 2-vertex model and introduce the dynamics proposed in [14, 17]. We will then study the evolution of the volume of the dual polyhedra to the 2-vertex graph. To do this, we will use the geometric quadrupole proposed in [27], which allows us to obtain an analytical expression that approximates the volume of the polyhedra and study its evolution.
III.1 Description and dynamics of the 2-vertex model
[…]
Figure 2: 2-vertex graph, labeled by $\alpha$ and $\beta$, with $N$ edges, and their respective spinors.
The simplicity of this model has made it possible to implement a non-trivial dynamics. With this aim, we consider the SU(2)-invariant observables [15, 14] (at the vertex $\alpha$)
@@ III.1 Description and dynamics of the 2-vertex model
$\lambda\in\mathbb{R},\gamma\in\mathbb{C}$ are constants with units of inverse length333 Note that $C$ is a Hamiltonian constraint that would come from integrating the Hamiltonian density from LQG over a fiducial spatial volume and therefore the coupling constants carry this information., and the lapse $\mathcal{M}$ is a Lagrange multiplier that enforces the Hamiltonian constraint $C=0$. This Hamiltonian is constructed to the lowest order in the operators (11) (treating both vertices equally). It is SU(2)-invariant, i.e., $\{H,\vec{\mathcal{X}}\}=\{H,\vec{\mathcal{Y}}\}=0$, where $\vec{\mathcal{X}}$ and $\vec{\mathcal{Y}}$ are the closure vector (10) at the vertices $\alpha$ and $\beta$ respectively. Furthermore, it satisfies the matching constraint, so it is U(1)-invariant on every edge, i.e. $\{H,\mathcal{C}_{i}\}=0$. For simplicity, the coupling constants $\lambda$ and $\gamma$ are taken edge independent but, in principle, they could also be different on each edge [14, 17]. Nevertheless, the symmetry reduction to the $\text{U}(N)$-invariant sector will impose edge independence on the constants [14], as it is already assumed in (12).
@@ III.2 Volume and quadrupole of a surface
Minkowski’s theorem [22] ensures the existence of a polyhedron associated with a set of vectors satisfying the closure constraint, but it does not give a prescription for the construction of the polyhedron. Although the polyhedron-reconstruction algorithm has recently been presented in [25, 26] and used in [17] in order to compute the volume numerically, there is no analytical formula for the volume of a polyhedron with more than $4$ faces in terms of the normal vectors to its faces. The tetrahedron ($N=4$) is an exception, and the volume squared is given by [27]:
@@
This quadrupole gives us basic information about the shape of the surface: its eigenvalues have a non-trivial relation with the principal radii of the ellipsoid that best approximates the surface in question, i.e. the ellipsoid with the same quadrupole [27].
Given that $X_{k}$ represents the area of the $k$-th face, the determinant of the geometrical quadrupole has an interpretation as a volume squared. Previous papers have explored the use of this determinant as an alternative to the volume squared of the polyhedron [27, 17], showing that both the determinant of the quadrupole and the volume squared (numerically calculated using the polyhedron-reconstruction algorithm presented in [25, 26]), although they do not coincide in value, have the same behavior in terms of their growth and extrema [17]. Therefore, the determinant of the quadrupole may be a good tool to study the evolution of the volume, for which we define the following function:
@@ III.3 Evolution of the volume
$\dot{\tilde{V}}\mathrel{\mathop{\ordinarycolon}}=\{H,\tilde{V}\}=\frac{\pi}{3\tilde{V}}\varepsilon^{IJK}\varepsilon^{LMN}\dot{T}_{IL}T_{JM}T_{KN}$ (16)
@@
where the dot denotes derivative with respect to the evolution parameter $t$ associated with the Hamiltonian $H$, and we have introduced the following (non gauge-invariant) functions:
@@
We observe in the analytical expression for the evolution of the volume (16) that a degenerate polyhedron (i.e. with all its faces parallel to each other) has null volume throughout the entire evolution. Similarly, we can obtain the evolution of the polyhedron associated with the vertex $\beta$ by substituting in equation (16) the functions and vectors defined on $\alpha$ by the corresponding ones defined on $\beta$ and vice versa.
@@
IV U($N$)-invariant sector of the 2-vertex model
The 2-vertex model presents an invariant sector under a global U$(N)$ symmetry (whose generators involve operators acting on both vertices) [15, 14]. One of the main results in this sector is that the equations of motion of the total area of the polyhedra are mathematically analogous to the equations of motion of volume in LQC [15]. This has lead to an interpretation of the 2-vertex model as a cosmological model. In this section we will review the U($N$) symmetry and we will study the dynamics of the 2-vertex model (fully described with spinors) on the U($N$)-invariant sector.
@@ IV.1 U($N$) symmetry
Indeed, we observe that the matching constraint associated with the edge $k$ is simply the $k$-th element $\mathcal{C}_{k}=\mathcal{E}_{kk}$ of the diagonal of $\mathcal{E}_{ij}$.
The generators $\mathcal{E}_{ij}$ are constants of motion, given that $\{\mathcal{E}_{ij},H\}=0$. Conversely, the Hamiltonian of the system is invariant under the U($N$) transformation generated by $\mathcal{E}_{ij}$. These transformations will be symmetries of the system only if the constraints (in this case the matching and closure constraints) are invariant under the action of $\mathcal{E}_{ij}$. If not, applying the U($N$) transformation generated by $\mathcal{E}_{ij}$ on configurations satisfying the constraints would lead to configurations not satisfying them. A simple calculation shows that, while the closure constraint is invariant, $\{\mathcal{E}_{ij},\vec{\mathcal{X}}\}=0$, the action of $\mathcal{E}_{ij}$ on the matching constraint (9) is non-trivial,

### v3

# From loop quantum gravity to cosmology: the two-vertex model

@@ Abstract
We study the notion of volume and its dynamics in the loop-quantum-gravity truncation known as the two-vertex model. We also show that its U$(N)$-symmetry reduction provides the old effective dynamics of loop quantum cosmology with an arbitrary perfect barotropic fluid content. A suitable modification of the Poisson bracket structure of the U($N$)-symmetric model leads to the loop quantum cosmology improved dynamics.
@@ I Introduction
The study of simple (truncated) models within the full theory has proved to be extremely useful. The truncation to a fixed graph is described by the corresponding holonomy-flux phase space on the graph and its quantization leads to the spin networks with support on that graph. The simplest nontrivial graphs are the two-vertex, $N$-edge graphs (the so-called two-vertex model, for brevity) [13, 14, 15, 16, 17, 18]. The two-vertex model can be interpreted as two polyhedra with identical areas but, in general, with different volume. The classical dynamics of these polyhedra has been studied analytically in a symmetry-reduced sector (by a global $\text{U}(N)$ symmetry) [14] and, more recently, numerically in the general case [17]. In both situations a cosmological behavior is observed, presenting oscillatory and divergent regimes that avoid any singularities. Besides, in [18] it was shown that, in the low-curvature regime, the $\text{U}(N)$-reduced two-vertex dynamics is related to the classical dynamics of a Friedmann-Lemaître-Robertson-Walker (FLRW) universe with a cosmological constant.
We will describe the two-vertex model using the spinorial formalism for LQG [19, 14, 20, 21], a description of the phase space of each edge using two spinors, one for each vertex of the edge. These spinors satisfy two constraints: the matching constraint—which ensures that there is only one SU(2) irreducible representation (one spin) on each edge—and the closure constraint—which ensures the SU(2) invariance of the theory. The spinorial formalism allows for a simple description of the model in terms of $2N$ spinors that, upon quantization, recovers the Hilbert space of LQG on the graph [19]. Therefore, we can study the classical dynamics of the model using the spinorial formalism in order to gain insight into full LQG, as it will give us a first contribution to the quantum dynamics of states peaked on classical trajectories. Moreover, using Minkowski’s theorem for convex polyhedra [22], the spinors provide a parametrization of the framework described by the twisted geometries [23, 24, 21]. In this framework, each vertex of the graph is associated with a polyhedron and the edges indicate adjacent faces with the same area but, in general, different shapes. In this way, the classical counterpart (parametrized by the spinors) associated with a spin network gives us an unconventional discretization of space in which the polyhedra do not fit, i.e., they are twisted with respect to each other.
@@
The $\text{U}(N)$-symmetric sector of the two-vertex model contains just one degree of freedom corresponding to the twist angle and its canonical conjugate momentum, i.e., the total area of the polyhedra. Comparing the evolution of the quadrupole approximation to the volume and the total area, we find that they precisely follow the evolution one expects from a homogeneous and isotropic evolution. We also find that a generalization of the two-vertex Hamiltonian constraint in the form of the addition of a function of the area (that preserves the symmetries of the reduced model) effectively takes into account matter content and, furthermore, fully determines the equation of state.
Finally, we show that the two-vertex model is indeed an LQG truncation that effectively describes the LQC old dynamics [28, 29] with arbitrary barotropic perfect fluids. Furthermore, a suitable modification of the Poisson bracket structure of the U($N$)-symmetric model leads to the LQC improved dynamics [9].
The paper is structured as follows. In Sec. II we introduce the spinorial formalism and the twisted geometries. We will use both of them in Sec. III to describe the two-vertex model and its dynamics. We will also use the notion of quadrupole moment of a discrete surface to study analytically the volume and its evolution. In Sec. IV we introduce the global $\text{U}(N)$ symmetry and study the dynamics on the reduced sector. In Sec. V, we show that the two-vertex model directly provides the FRLW cosmology and the LQC effective dynamics. We finally summarize and conclude in Sec. VI.
Index notation. $a,b\ldots=1,2,3$ are spatial indices; $i,j\ldots=1,2...,N$ denote the edges of the two-vertex graph; $I,J\ldots=1,2,3$ are $\mathfrak{su}(2)$ indices; and $A,B=0,1$ are spinorial indices.
@@ II.1 Spinorial formalism for LQG
This parametrization of $T^{*}\mathrm{SU(2)}$ in terms of two spinors is the classical analog to Schwinger’s representation of SU(2) by two decoupled harmonic oscillators  [30]. According to this representation, by canonically quantizing the spinors with the commutation relations (3), we will obtain two harmonic oscillators on each edge, describing the SU(2) irreducible representation associated with such edge [30, 15, 19].
@@
$g_{i}:=\frac{|z_{i}^{t}]\bra{z_{i}^{s}}-\ket{z_{i}^{t}}[z_{i}^{s}|}{\sqrt{\braket{z_{i}^{t}}{z_{i}^{t}}\braket{z_{i}^{s}}{z_{i}^{s}}}},$ (5)
$\vec{X}^{s,t}_{i}\!:=\!\frac{1}{2}\!\bra{z_{i}^{s,t}}\vec{\sigma}\ket{z_{i}^{s,t}},\quad\!\!X^{s,t}_{i}:=|\vec{X}^{s,t}_{i}|\!=\!\frac{1}{2}\!\braket{z_{i}^{s,t}}{z_{i}^{s,t}}.$ (6)
@@
The Barbero-Immirzi parameter $\beta$ divides the flux and nontrivially modifies the holonomy, in such a way that the Poisson bracket is invariant under the transformation (4) as happens in LQG. The action of $g_{i}$ associates the (normalized) spinors on one vertex with the dual (normalized) spinors on the other vertex:
@@ II.2 Twisted geometries
The closure constraint (10) on noncoplanar vectors is precisely the condition required by Minkowski’s theorem [22]. Therefore, these vectors define a unique polyhedron at each vertex, whose faces are orthogonal to the vectors and its areas equal their norms. The faces associated with the same edge but different vertex are understood as adjacent, and the matching constraint (9) causes both faces on each edge to have the same area. However, there is no restriction on the shape of the faces. The faces associated with the same edge do not necessarily have to fit together, giving a notion of twist between adjacent faces. The twist on each edge is parametrized by the twist angle $\xi_{i}\in[-\pi,\pi]$. Consequently, a graph will describe a discretization of space, in which the polyhedra that share an edge will have an adjacent face with the same area but different shapes. For this reason this formalism is known as twisted geometries [23, 24]. According to this formalism, the phase space corresponding to each edge $i$ connecting two vertices is described by $(\hat{X}^{s}_{i},\hat{X}^{t}_{i},X_{i},\xi_{i})$, where $\hat{X}^{s}_{i}$ and $\hat{X}^{t}_{i}$ are the normal vectors to the source and target faces corresponding to the edge $i$; $X_{i}:=X_{i}^{s}=X_{i}^{t}$ is the area of the faces associated with the edge $i$; and $\xi_{i}$ is the twist angle of that edge. See Fig. (1) for a graphical representation of the polyhedra associated with a given graph.
@@
III Two-vertex model
We will now apply the formalisms explained in the previous section to describe the two-vertex model and introduce the dynamics proposed in [14, 17]. We will then study the evolution of the volume of the dual polyhedra to the two-vertex graph. To do this, we will use the geometric quadrupole proposed in [27], which allows us to obtain an analytical expression that approximates the volume of the polyhedra and study its evolution.
III.1 Description and dynamics of the two-vertex model
Let us consider the two-vertex model. We will label the vertices [^2] Note that the concept of source and target vertices depends on the edge, while the labels $\alpha$ and $\beta$ characterize concrete vertices. This notation is more convenient for the two-vertex model. by $\alpha$ and $\beta$ (see Fig. 2[17, 15, 14]. Following the standard notation in the literature, we will denote spinors at vertex $\alpha$ as |$\ket{z_{i}}:=\ket{z^{\alpha}_{i}}$, spinors at vertex $\beta$ as |$\ket{w_{i}}:=\ket{z^{\beta}_{i}}$, and the vectors at each vertex as $\vec{X}_{i}:=\vec{X}_{i}^{\alpha}$ and $\vec{Y}_{i}:=\vec{X}_{i}^{\beta}$.
Figure 2: Two-vertex graph, labeled by $\alpha$ and $\beta$, with $N$ edges, and their respective spinors.
The simplicity of this model has made it possible to implement a nontrivial dynamics. With this aim, we consider the SU(2)-invariant observables [15, 14] (at the vertex $\alpha$)
@@ III.1 Description and dynamics of the two-vertex model
$\lambda\in\mathbb{R},\gamma\in\mathbb{C}$ are constants with units of inverse length333 Note that $C$ is a Hamiltonian constraint that would come from integrating the Hamiltonian density from LQG over a fiducial spatial volume and therefore the coupling constants carry this information., and the lapse $\mathcal{M}$ is a Lagrange multiplier that enforces the Hamiltonian constraint $C=0$. This Hamiltonian is constructed to the lowest order in the operators (11) (treating both vertices equally). It is SU(2)-invariant, i.e., $\{H,\vec{\mathcal{X}}\}=\{H,\vec{\mathcal{Y}}\}=0$, where $\vec{\mathcal{X}}$ and $\vec{\mathcal{Y}}$ are the closure vector (10) at the vertices $\alpha$ and $\beta$, respectively. Furthermore, it satisfies the matching constraint, so it is U(1) invariant on every edge, i.e. $\{H,\mathcal{C}_{i}\}=0$. For simplicity, the coupling constants $\lambda$ and $\gamma$ are taken edge independent but, in principle, they could also be different on each edge [14, 17]. Nevertheless, the symmetry reduction to the $\text{U}(N)$-invariant sector will impose edge independence on the constants [14], as it is already assumed in (12).
@@ III.2 Volume and quadrupole of a surface
Minkowski’s theorem [22] ensures the existence of a polyhedron associated with a set of vectors satisfying the closure constraint, but it does not give a prescription for the construction of the polyhedron. Although the polyhedron-reconstruction algorithm has recently been presented in [25, 26] and used in [17] in order to compute the volume numerically, there is no analytical formula for the volume of a polyhedron with more than four faces in terms of the normal vectors to its faces. The tetrahedron ($N=4$) is an exception, and the volume squared is given by [27]:
@@
This quadrupole gives us basic information about the shape of the surface: its eigenvalues have a nontrivial relation with the principal radii of the ellipsoid that best approximates the surface in question, i.e. the ellipsoid with the same quadrupole [27].
Given that $X_{k}$ represents the area of the $k$th face, the determinant of the geometrical quadrupole has an interpretation as a volume squared. Previous papers have explored the use of this determinant as an alternative to the volume squared of the polyhedron [27, 17], showing that both the determinant of the quadrupole and the volume squared (numerically calculated using the polyhedron-reconstruction algorithm presented in [25, 26]), although they do not coincide in value, have the same behavior in terms of their growth and extrema [17]. Therefore, the determinant of the quadrupole may be a good tool to study the evolution of the volume, for which we define the following function:
@@ III.3 Evolution of the volume
$\dot{\tilde{V}}:=\{H,\tilde{V}\}=\frac{\pi}{3\tilde{V}}\varepsilon^{IJK}\varepsilon^{LMN}\dot{T}_{IL}T_{JM}T_{KN}$ (16)
@@
where the dot denotes derivative with respect to the evolution parameter $t$ associated with the Hamiltonian $H$, and we have introduced the following (non-gauge-invariant) functions:
@@
We observe in the analytical expression for the evolution of the volume (16) that a degenerate polyhedron (i.e. with all its faces parallel to each other) has null volume throughout the entire evolution. Similarly, we can obtain the evolution of the polyhedron associated with the vertex $\beta$ by substituting in Eq. (16) the functions and vectors defined on $\alpha$ by the corresponding ones defined on $\beta$ and vice versa.
@@
IV U($N$)-invariant sector of the two-vertex model
The two-vertex model presents an invariant sector under a global U$(N)$ symmetry (whose generators involve operators acting on both vertices) [15, 14]. One of the main results in this sector is that the equations of motion of the total area of the polyhedra are mathematically analogous to the equations of motion of volume in LQC [15]. This has lead to an interpretation of the two-vertex model as a cosmological model. In this section we will review the U($N$) symmetry and we will study the dynamics of the two-vertex model (fully described with spinors) on the U($N$)-invariant sector.
@@ IV.1 U($N$) symmetry
Indeed, we observe that the matching constraint associated with the edge $k$ is simply the $k$th element $\mathcal{C}_{k}=\mathcal{E}_{kk}$ of the diagonal of $\mathcal{E}_{ij}$.
The generators $\mathcal{E}_{ij}$ are constants of motion, given that $\{\mathcal{E}_{ij},H\}=0$. Conversely, the Hamiltonian of the system is invariant under the U($N$) transformation generated by $\mathcal{E}_{ij}$. These transformations will be symmetries of the system only if the constraints (in this case the matching and closure constraints) are invariant under the action of $\mathcal{E}_{ij}$. If not, applying the U($N$) transformation generated by $\mathcal{E}_{ij}$ on configurations satisfying the constraints would lead to configurations not satisfying them. A simple calculation shows that, while the closure constraint is invariant, $\{\mathcal{E}_{ij},\vec{\mathcal{X}}\}=0$, the action of $\mathcal{E}_{ij}$ on the matching constraint (9) is nontrivial,